### v1

@@ 1.3 Main Research Gaps and Objectives
  • To ensure that the variational optimization of our ATTBHFA-Net can be performed in an analogous footing to that of the other related methods (e.g., VAE [47], WVAE [31]), we derived a mathematical relation linking our modified ELBO (which would be known as the EVLBO) and the Bhattacharyya coefficient. The corresponding relationship between the modified ELBO and the Hellinger distance was made in [40]. These relations are and have been shown to be mathematically consistent like that of the relations obtained from other variational-based works.
@@ 2.1 On Few-Shot Learning Review
Most of the FSL-based approaches discussed in the following can be categorized into metric-based learning approaches, in which a distance function over the data samples is learned [50]. The objective is to learn an embedding function in the embedded space given a distance function, with each function characterized by the relevant parameters, which is computed by the neural network. The work by Bateni et al. [33] highlighted the importance of selecting the distance metric, which has been largely unfocused on previously due to the common perceptions that non-linear mappings via neural networks should presumably adapt to any kind of metric, and hence making any selected distance choices inconsequential.
@@ 4.1.2 Bhattarcharyya Distance
The Bhattarcharyya distance $D_{BD}$ in our ATTBHFA-Net is mathematically depicted as
$D_{BD}=-\textrm{ln}\left(\int\sqrt{(p_{\theta}(z|\mathcal{T})q_{\phi}(z|\mathcal{S}))}dz\right).$ (5)
@@
To ensure that the variation optimization can be done on a similar footing as that of the KL divergence and Wasserstein distance, we can derived a modified Evidence Variational Lower Bound (EVLBO) which is related to the Bhattarcharyya coefficient in a similar manner as that of Lee et al. [40]. The mathematical derivation is detailed in the appendix section, but we denote the final mathematical result as follows:
$EVLBO=\int\textrm{log}\left(p_{\theta}(\mathcal{T})\right)dz+2\left(\textrm{log}(BC)-\int\textrm{log}\left(q_{\phi}(z|\mathcal{S})\right)dz\right),$ (6)
where $\int\textrm{log}(p_{\theta}(\mathcal{T}))dz$ is the evidence term and $EVLBO$ can be expressed as $\frac{ELBO}{q_{\phi}(z|\mathcal{S})}$. Therefore, EVLBO is related to ELBO divided by the posterior distribution $q_{\phi}{(z|\mathcal{S})}$, and its value is dependent on the latter, hence its name. It can be shown that if $p_{\theta}(z|\mathcal{T})=q_{\phi}(z|\mathcal{S})$, the $EVLBO$ is equivalent to the evidence term, in agreement with previous related work that associates the respective ELBO with the KL divergence ([29], [66]), the Wasserstein distance [31] and the Hellinger distance [40].
@@ 4.1.3 Hellinger Distance
As derived in [40] and [41], $D_{H}^{2}$ can be related to the modified ELBO as
$\begin{split}ELBO^{\prime}=\int\textrm{log}(p_{\theta}(\mathcal{T}))dz+\textrm{log}\left(1-D^{2}_{H}\right)^{2},\end{split}$ (8)
where the new ELBO term, $ELBO^{\prime}$, is expressed as $\frac{ELBO}{q_{\phi}(z|\mathcal{S})}$.
Together, we can express the optimization objective of the ATTBHFA-Net (to maximize the $EVLBO$ and $ELBO^{\prime}$ (equivalent to minimizing the Bhattarcharyya coefficient and Hellinger distance) as
$\underset{\phi,\theta}{\arg\max}\sum_{x\in\mathcal{\mathcal{T},\mathcal{S}}}EVLBO+\underset{\phi,\theta}{\arg\max}\sum_{x\in\mathcal{T},\mathcal{S}}ELBO^{\prime}.$ (9)
@@ 7 Conclusions
Appendix A Derivation of the Modified ELBO (EVLBO) for Bhattarcharyya Coefficient
@@
In this appendix section, we provide a derivation for the modified ELBO (EVLBO) that relates to the Bhattarcharyya coefficient $BC$. We start with the derivation of the original ELBO as
@@
$ELBO=\int q_{\phi}(z|\mathcal{S})\textrm{log}\left(\frac{p_{\theta}(\mathcal{T},z)}{q_{\phi}(z|\mathcal{S})}\right)dz$ (19)
and that the Bhattarcharyya coefficient is written as
$BC=\int\sqrt{p_{\theta}(z|\mathcal{T})q_{\phi}(z|\mathcal{S})}dz.$ (20)
Taking the logarithm of each term on both sides, we get
$\textrm{log}(BC)=\frac{1}{2}\textrm{log}\left(\int p_{\theta}(z|\mathcal{T})dz\right)+\frac{1}{2}\textrm{log}\left(\int q_{\phi}(z|\mathcal{S})dz\right).$ (21)
Since $p_{\theta}(\mathcal{T},z)=p_{\theta}(z|\mathcal{T})p_{\theta}(\mathcal{T})$, equation 21 can be written as
$\textrm{log}(BC)=\frac{1}{2}\textrm{log}\left(\int\frac{p_{\theta}(z,\mathcal{T})}{p_{\theta}(\mathcal{T})}dz\right)+\frac{1}{2}\textrm{log}\left(\int q_{\phi}(z|\mathcal{S})dz\right)$ (22)
which we can cast $\textrm{log}\left(\int\frac{p_{\theta}(z,\mathcal{T})}{p_{\theta}(\mathcal{T})}dz\right)$ as $\textrm{log}\left(\int p_{\theta}(z,\mathcal{T})dz\right)-\textrm{log}\left(\int p_{\theta}(\mathcal{T})dz\right)$, and bring the second term to the left side of equation 22 to obtain
$\textrm{log}(BC)+\frac{1}{2}\textrm{log}\left(\int p_{\theta}(\mathcal{T})dz\right)=\frac{1}{2}\textrm{log}\left(\int p_{\theta}(z,\mathcal{T})dz\right)+\frac{1}{2}\textrm{log}\left(\int q_{\phi}(z|\mathcal{S})dz\right),$ (23)
where we note that the term $\int p_{\theta}(\mathcal{T})dz$ is the evidence term. We can insert $\frac{q_{\phi}(z|\mathcal{S})}{q_{\phi}(z|\mathcal{S})}$ in $\frac{1}{2}log\left(\int p_{\theta}(z,\mathcal{T})dz\right)$ to yield
$\begin{split}\textrm{log}(BC)+\frac{1}{2}\textrm{log}\left(\int p_{\theta}(\mathcal{T})dz\right)=\frac{1}{2}\textrm{log}\left(\int p_{\theta}(z,\mathcal{T})\frac{q_{\phi}(z|\mathcal{S})}{q_{\phi}(z|\mathcal{S})}dz\right)+\frac{1}{2}\textrm{log}\left(\int q_{\phi}(z|\mathcal{S})dz\right)\\ =\frac{1}{2}\textrm{log}\left(\int\frac{p_{\theta}(z,\mathcal{T})}{q_{\phi}(z|\mathcal{S})}dz\right)+\frac{1}{2}\textrm{log}\left(\int q_{\phi}(z|\mathcal{S})dz\right)+\frac{1}{2}\textrm{log}\left(\int q_{\phi}(z|\mathcal{S})dz\right)\\ =\frac{1}{2}\textrm{log}\left(\int\frac{p_{\theta}(z,\mathcal{T})}{q_{\phi}(z|\mathcal{S})}dz\right)+\textrm{log}\left(\int q_{\phi}(z|\mathcal{S})dz\right).\end{split}$ (24)
In all the terms above, the logarithm can be brought into the integral, and hence we can rewrite equation 24 as
$\textrm{log}(BC)+\frac{1}{2}\int\textrm{log}\left(p_{\theta}(\mathcal{T})\right)dz=\frac{1}{2}\int\textrm{log}\left(\frac{p_{\theta}(z,\mathcal{T})}{q_{\phi}(z|\mathcal{S})}\right)dz+\int\textrm{log}\left(q_{\phi}(z|\mathcal{S})\right)dz.$ (25)
Looking back to the definition of ELBO, we can make $log\left(\frac{p_{\theta}(z,\mathcal{T})}{q_{\phi}(z|\mathcal{S})}\right)$ as the subject of the equation and obtain
$\frac{1}{q_{\phi}(z|\mathcal{S})}\frac{d(ELBO)}{dz}=\textrm{log}\left(\frac{p_{\theta}(z,\mathcal{T})}{q_{\phi}(z|\mathcal{S})}\right)$ (26)
which implies
@@
$\begin{split}\textrm{log}(BC)+\frac{1}{2}\int\textrm{log}\left(p_{\theta}(\mathcal{T})\right)dz=\frac{1}{2}\int\frac{1}{q_{\phi}(z|\mathcal{S})}\frac{d(ELBO)}{dz}dz+\int\textrm{log}\left(q_{\phi}(z|\mathcal{S})\right)dz\\ =\frac{1}{2}\frac{ELBO}{q_{\phi}(z|\mathcal{S})}+\int\textrm{log}\left(q_{\phi}(z|\mathcal{S})\right)dz.\end{split}$ (27)
Bringing $\int\textrm{log}\left(q_{\phi}(z|\mathcal{S})\right)dz$ to the left-hand side of equation 27 and bringing out the factor of 2 on both sides,
$2\textrm{log}(BC)+\int\textrm{log}\left(p_{\theta}(\mathcal{T})\right)dz-2\int\textrm{log}\left(q_{\phi}(z|\mathcal{S})\right)dz=\frac{ELBO}{q_{\phi}(z|\mathcal{S})},$ (28)
which we defined the modified ELBO (EVLBO) as $EVLBO=\frac{ELBO}{q_{\phi}(z|\mathcal{S})}$, we finally obtain
$2\textrm{log}(BC)+\int\textrm{log}\left(p_{\theta}(\mathcal{T})\right)dz-2\int\textrm{log}\left(q_{\phi}(z|\mathcal{S})\right)dz=EVLBO.$ (29)
Therefore, the EVLBO is associated with the Bhattarcharyya coefficient via
$\begin{split}EVLBO=\int\textrm{log}\left(p_{\theta}(\mathcal{T})\right)dz+2\textrm{log}(BC)-2\int\textrm{log}\left(q_{\phi}(z|\mathcal{S})\right)dz\\ =\int\textrm{log}\left(p_{\theta}(\mathcal{T})\right)dz+2\left(\textrm{log}(BC)-\int\textrm{log}\left(q_{\phi}(z|\mathcal{S})\right)dz\right).\end{split}$ (30)
The term in the parenthesis of the right hand side of equation 30 becomes $\frac{1}{2}\int\textrm{log}\left(p_{\theta}(z|\mathcal{T})\right)dz-\frac{1}{2}\int\textrm{log}\left(q_{\phi}(z|\mathcal{S})\right)dz$, since $\textrm{log}(BC)=\frac{1}{2}\int\textrm{log}\left(p_{\theta}(z|\mathcal{T})\right)dz+\frac{1}{2}\int\textrm{log}\left(q_{\phi}(z|\mathcal{S})\right)dz$. From here, we can easily see that if $p_{\theta}(z|\mathcal{T})=q_{\phi}(z|\mathcal{S})$, the term in the parenthesis becomes 0, and hence the evidence term equates to the EVLBO, in agreement with the earlier statement in the paper.
In summary, we can expressed the optimization objective of the Bhattacharyya distance component of the ATTBHFA-Net (to maximize the EVLBO) as
$\underset{\phi,\theta}{\arg\max}\sum_{x\in\mathcal{\mathcal{T},\mathcal{S}}}EVLBO.$ (31)

### v2

@@ 1.3 Main Research Gaps and Objectives
  • To ensure that the variational optimization of our ATTBHFA-Net can be performed in an analogous footing to that of the other related methods (e.g., VAE [47], WVAE [31]), we derived a mathematical relation linking our modified ELBO (which would be known as the $ELBO^{*}$) and the Bhattacharyya coefficient. The corresponding relationship between the modified ELBO and the Hellinger distance was made in [40]. These relations are and have been shown to be mathematically consistent like that of the relations obtained from other variational-based works.
@@ 2.1 On Few-Shot Learning Review
Most of the FSL-based approaches discussed in the following can be categorized into metric-based learning approaches, in which a distanceEV function over the data samples is learned [50]. The objective is to learn an embedding function in the embedded space given a distance function, with each function characterized by the relevant parameters, which is computed by the neural network. The work by Bateni et al. [33] highlighted the importance of selecting the distance metric, which has been largely unfocused on previously due to the common perceptions that non-linear mappings via neural networks should presumably adapt to any kind of metric, and hence making any selected distance choices inconsequential.
@@ 4.1.2 Bhattarcharyya Distance
The Bhattarcharyya distance $D_{B}$ in our ATTBHFA-Net is mathematically depicted as
$D_{B}=-\textrm{ln}\left(\int\sqrt{(p_{\theta}(z|\mathcal{T})q_{\phi}(z|\mathcal{S})}dz\right).$ (5)
@@
To ensure that the variation optimization can be done on a similar footing as that of the KL divergence and Wasserstein distance, we can derived a modified Evidence Lower Bound ($ELBO^{*}$) which is related to the Bhattarcharyya distance (and hence coefficient) in a similar manner as that of Lee et al. [40]. The mathematical derivation is detailed in the appendix section, but we denote the final mathematical result as follows:
$\begin{split}ELBO^{*}=\int q_{\phi}(z|\mathcal{S})\textrm{ln}\left(p_{\theta}(z|\mathcal{T})\right)dz+2D_{B}\\ =\int q_{\phi}(z|\mathcal{S})\textrm{ln}\left(p_{\theta}(z|\mathcal{T})\right)dz-2ln(BC)\end{split}$ (6)
where $\int q_{\phi}(z|\mathcal{S})\textrm{ln}\left(p_{\theta}(\mathcal{T}|z)\right)dz$ is the reconstruction term and $ELBO^{*}$ can be expressed as $\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-ELBO$. It can be shown that if $p_{\theta}(z|\mathcal{T})=q_{\phi}(z|\mathcal{S})$, $ELBO^{*}$ is equivalent to the reconstruction term, in agreement with previous related work that associates the respective ELBO with the KL divergence ([29], [66]), the Wasserstein distance [31] and the Hellinger distance [40].
@@ 4.1.3 Hellinger Distance
As derived from [40] and [41], $D_{H}^{2}$ can be related to the modified ELBO as
$\begin{split}ELBO^{\prime}=\int\textrm{log}(p_{\theta}(\mathcal{T}))dz+log\left(1-D^{2}_{H}\right)^{2},\end{split}$ (8)
where $ELBO^{\prime}$ is the new ELBO term expressed as $\frac{ELBO}{q_{\phi}(z|\mathcal{S})}$ in the original paper, and $\int log(p_{\theta}(\mathcal{T}))dz$ is the evidence term. However, in the appendix we will derive the aforementioned relationship again from another mathematical angle, and arrived at
$\begin{split}ELBO^{\prime\prime}=-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz+log(1-D_{H}^{2})^{2}\end{split}$ (9)
where $ELBO^{\prime\prime}=ELBO-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$. We note that
$\begin{split}\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz=\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T})p_{\theta}(\mathcal{T}))dz=\\ \int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz+\int q_{\phi}(z|\mathcal{S})p_{\theta}(\mathcal{T})dz.\end{split}$ (10)
So we can rewrite $\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz$ as $\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz+\int q_{\phi}(z|\mathcal{S})p_{\theta}(\mathcal{T})dz$. Move the term $\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$ to the left-hand side of equation 9 and the term $\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$ cancels out in $ELBO^{\prime\prime}$. By dividing both sides by $q_{\phi}(z|\mathcal{S})$, we recovered equation 8.
Together, we can express the optimization objective of the ATTBHFA-Net (to maximize $ELBO^{*}$ and $ELBO^{\prime\prime}$ (equivalent to minimizing the Bhattarcharyya coefficient and Hellinger distance) as
$\underset{\phi,\theta}{\arg\max}\sum_{x\in\mathcal{\mathcal{T},\mathcal{S}}}ELBO^{*}+\underset{\phi,\theta}{\arg\max}\sum_{x\in\mathcal{T},\mathcal{S}}ELBO^{\prime\prime}.$ (11)
@@ 7 Conclusions
Appendix A Derivation of the Modified ELBO ($ELBO^{*}$) for Bhattarcharyya Distance (and Coefficient)
@@
In this appendix section, we provide a derivation for the modified ELBO ($ELBO^{*}$) that relates to the Bhattarcharyya distance $D_{B}$ (and hence the coefficient $BC$). We start with the definition of $BC$ as
$BC=\int\sqrt{q_{\phi}(z|\mathcal{S})p_{\theta}(z|\mathcal{T})}dz$ (21)
The derivation works best if we consider the Bhattarcharryya distance $D_{B}$ directly, which is defined as the negative of the natural logarithm of $BC$ as $D_{B}=-ln(BC)$. From this we take the derivative with respect to $z$ on both sides to get
$\frac{d}{dz}\left(BC\right)=\sqrt{q_{\phi}(z|\mathcal{S})p_{\theta}(z|\mathcal{T})}.$ (22)
Taking natural logarithm on both sides and expanding out the right-hand side, we get
$ln\left(\frac{d}{dz}\left(e^{-D_{B}}\right)\right)=\frac{1}{2}ln(q_{\phi}(z|\mathcal{S}))+\frac{1}{2}ln(p_{\theta}(z|\mathcal{T})).$ (23)
We now use the property that $\int q_{\phi}(z|\mathcal{S})dz=1$ and introduced it on the left-hand side of the equation to get
$ln\left(\frac{d}{dz}\left(e^{-D_{B}}\int q_{\phi}(z|\mathcal{S})dz\right)\right)=\frac{1}{2}ln(q_{\phi}(z|\mathcal{S}))+\frac{1}{2}ln(p_{\theta}(z|\mathcal{T})),$ (24)
where in general, $D_{B}$ is dependent on z, but in the VAE context, the posterior and the prior distributions have means and variances independent of $z$, so we can bring the $e^{-D_{B}}$ term in the integral to obtain
$ln\left(\frac{d}{dz}\int q_{\phi}(z|\mathcal{S})e^{-D_{B}}dz\right)=\frac{1}{2}ln(q_{\phi}(z|\mathcal{S}))+\frac{1}{2}ln(p_{\theta}(z|\mathcal{T})),$ (25)
$ln\left(q_{\phi}(z|\mathcal{S})e^{-D_{B}}\right)=\frac{1}{2}ln(q_{\phi}(z|\mathcal{S}))+\frac{1}{2}ln(p_{\theta}(z|\mathcal{T})).$ (26)
Expanding out the left-hand side, taking note that $ln(e^{-D_{B}})=-D_{B}$,
$ln\left(q_{\phi}(z|\mathcal{S})\right)-D_{B}=\frac{1}{2}ln(q_{\phi}(z|\mathcal{S}))+\frac{1}{2}ln(p_{\theta}(z|\mathcal{T}))$ (27)
and multiplying both sides by $q_{\phi}(z|\mathcal{S})$, followed by integrating both sides with respect to dz, we get
$\begin{split}\int q_{\phi}(z|\mathcal{S})ln\left(q_{\phi}(z|\mathcal{S})\right)dz-\int D_{B}q_{\phi}(z|\mathcal{S})dz=\\ \frac{1}{2}\int q_{\phi}(z|\mathcal{S})ln(q_{\phi}(z|\mathcal{S}))dz+\frac{1}{2}\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz.\end{split}$ (28)
We note that $\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz$ is nothing but the reconstruction term. Rearranging,
$\frac{1}{2}\int q_{\phi}(z|\mathcal{S})ln\left(q_{\phi}(z|\mathcal{S})\right)dz-\frac{1}{2}\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz=\int D_{B}q_{\phi}(z|\mathcal{S})dz,$ (29)
$\int q_{\phi}(z|\mathcal{S})ln\left(q_{\phi}(z|\mathcal{S})\right)dz-\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz=2\int D_{B}q_{\phi}(z|\mathcal{S})dz.$ (30)
Inserting $\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$ on the left-hand side while shifting the reconstruction term to the right,
$\begin{split}\int q_{\phi}(z|\mathcal{S})ln\left(q_{\phi}(z|\mathcal{S})\right)dz+\left(\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz\right)=\\ \int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz+2\int D_{B}q_{\phi}(z|\mathcal{S})dz,\end{split}$ (31)
@@
and we can obtain the negative of the ELBO as $ELBO=\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz=-\int q_{\phi}(z|\mathcal{S})ln\left(\frac{p_{\theta}(z|\mathcal{S},\mathcal{T})}{q_{\phi}(z|\mathcal{S})}\right)dz$,
$\begin{split}\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-ELBO=\\ \int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz+2\int D_{B}q_{\phi}(z|\mathcal{S})dz.\end{split}$ (32)
We make use of a property that is useful (especially during the image reconstruction process): Since $D_{B}$ and thus $BC$ is measured via comparing the distribution $p_{\theta}(z|\mathcal{T})$ and $q_{\phi}(z|\mathcal{S})$, and in VAE the latter can be expressed in terms of the means and covariances of the distribution, all independent of $z$ as they depend on the original input images, we can bring $D_{B}$ out of the integral and utilize $\int q_{\phi}(z|\mathcal{S})=1$ again to get
$\begin{split}\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-ELBO=\\ \int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz+2D_{B}.\end{split}$ (33)
Finally, define the $ELBO^{*}$ as $\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-ELBO$, we arrived at
@@
$\begin{split}ELBO^{*}=\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz+2D_{B}=\int q_{\phi}(z|\mathcal{S})\textrm{ln}\left(p_{\theta}(z|\mathcal{T})\right)dz-2ln(BC)\end{split}$ (34)
which is equation 6.
Appendix B Another Derivation of the Modified ELBO ($ELBO^{\prime\prime}$) for the Hellinger Distance
In this appendix section, we provide another derivation for the modified ELBO ($ELBO^{\prime\prime}$) that relates to the square of the Hellinger distance $D_{H}^{2}$ (which has already been in [40] and [41]). We start with the definition of the $D_{H}^{2}$ as
$D_{H}^{2}=1-\int\sqrt{p_{\theta}(z|\mathcal{T})q_{\phi}(z|\mathcal{S})}dz.$ (35)
Switching the terms on each side to get
$\int\sqrt{p_{\theta}(z|\mathcal{T})q_{\phi}(z|\mathcal{S})}dz=1-D_{H}^{2}.$ (36)
Take derivative with respect to $z$ on both sides,
$\frac{d}{dz}\int\sqrt{p_{\theta}(z|\mathcal{T})q_{\phi}(z|\mathcal{S})}dz=\frac{d}{dz}(1-D_{H}^{2}),$ (37)
$\sqrt{p_{\theta}(z|\mathcal{T})q_{\phi}(z|\mathcal{S})}=\frac{d}{dz}(1-D_{H}^{2}),$ (38)
and taking logarithm on both sides,
$\frac{1}{2}log(p_{\theta}(z|\mathcal{T}))+\frac{1}{2}log(q_{\phi}(z|\mathcal{S}))=log\left(\frac{d}{dz}(1-D_{H}^{2})\right),$ (39)
$log(p_{\theta}(z|\mathcal{T}))+log(q_{\phi}(z|\mathcal{S}))=2log\left(\frac{d}{dz}(1-D_{H}^{2})\right).$ (40)
Now we apply an analogous argument as in the previous derivation: Since $D^{2}_{H}$ is measured via comparing the distribution $p_{\theta}(z|\mathcal{T})$ and $q_{\phi}(z|\mathcal{S})$, and in VAE the latter can be expressed in terms of the means and covariances of the distribution, all independent of $z$ as they depend on the original input images, $log\left(\frac{d}{dz}(1-D_{H}^{2})\right)$ is independent of $z$ and we can insert $\int q_{\phi}(z|\mathcal{S})dz=1$ in the parenthesis of $(1-D_{H}^{2})$ to get
$\begin{split}log(p_{\theta}(z|\mathcal{T}))+log(q_{\phi}(z|\mathcal{S}))=2log\left(\frac{d}{dz}(1-D_{H}^{2})\int q_{\phi}(z|\mathcal{S})dz\right)\\ =2log\left(\frac{d}{dz}\int(1-D_{H}^{2})q_{\phi}(z|\mathcal{S})dz\right)\\ =2log\left((1-D_{H}^{2})q_{\phi}(z|\mathcal{S})\right)\\ =2log(1-D_{H}^{2})+2log\left(q_{\phi}(z|\mathcal{S}\right).\end{split}$ (41)
Bringing $2log(q_{\phi}(z|\mathcal{S})$ to the left-hand side of the equation,
$\begin{split}log(p_{\theta}(z|\mathcal{T}))-log(q_{\phi}(z|\mathcal{S}))=2log(1-D_{H}^{2}).\end{split}$ (42)
Multiplying both sides by $q_{\phi}(z|\mathcal{S})$ and integrate them with respect to $z$,
$\begin{split}\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz-\int q_{\phi}(z|\mathcal{S})log(q_{\phi}(z|\mathcal{S}))dz=2\int q_{\phi}(z|\mathcal{S})log(1-D_{H}^{2})dz.\end{split}$ (43)
Use the property that $log(1-D_{H}^{2})$ is independent of $z$ to bring it out of the integral on the right:
$\begin{split}\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz-\int q_{\phi}(z|\mathcal{S})log(q_{\phi}(z|\mathcal{S}))dz=2\left(\int q_{\phi}(z|\mathcal{S})dz\right)log(1-D_{H}^{2})\\ =2log(1-D_{H}^{2}),\end{split}$ (44)
where in the last line we made use of $\int q_{\phi}(z|\mathcal{S})dz=1$ again. Recognizing that $\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz$ is the reconstruction term, we insert $\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$ on the left-hand side while shifting the reconstruction term to the right,
$\begin{split}\left(\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz\right)-\int q_{\phi}(z|\mathcal{S})log(q_{\phi}(z|\mathcal{S}))dz=\\ -\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz+2log(1-D_{H}^{2}),\end{split}$ (45)
$\begin{split}\left(\int q_{\phi}(z|\mathcal{S})log\left(\frac{p_{\theta}(z|\mathcal{S},\mathcal{T})}{q_{\phi}(z|\mathcal{S})}\right)dz-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz\right)=\\ -\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz+2log(1-D_{H}^{2}),\end{split}$ (46)
$\begin{split}ELBO-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz=-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz+2log(1-D_{H}^{2}),\end{split}$ (47)
and we finally arrived at
$\begin{split}ELBO^{\prime\prime}=-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz+log(1-D_{H}^{2})^{2},\end{split}$ (48)
which is equation 9. The modified $ELBO^{\prime\prime}$ reads $ELBO-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$. Some remarks are in order:
  • The reconstruction term in the relationship between $ELBO^{*}$ and $D_{B}$ involved a natural logarithm of $p_{\theta}(z|\mathcal{T})$, while the reconstruction term in the relationship between $ELBO^{\prime\prime}$ and $D^{2}_{H}$ involved a logarithm of base 10 of $p_{\theta}(z|\mathcal{T})$. This is because in our derivation of the relationship between $ELBO^{*}$ and $D_{B}$, we encountered a term of the form $ln(e^{-D_{B}})$, which if expressed in log base 10 could not be reduced to $D_{B}$.
  • The reconstruction term in the relationship between $ELBO^{*}$ and $D_{B}$ is positive on the right-hand side of the equation while the reconstruction term in the relationship between $ELBO^{\prime\prime}$ and $log(1-D_{H}^{2})^{2}$ is negative on the right-hand side of the equation.
  • The modified ELBO in equation 6 ($ELBO^{*}$) involved subtraction by the ELBO term while the modified ELBO in equation 9 $($$ELBO^{\prime\prime}$) involved subtraction directly from ELBO. This means that in the overall optimization objective as shown in equation 11, the original ELBO term cancels out and the modified ELBO summation in the maximization of the objective mainly involved the reconstruction term $\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$ and $\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$. This also implies that the summation involved the term containing the square of the Hellinger distance $log(1-D_{H}^{2})^{2}$ getting subtracted by twice of the Bhattarcharyya coefficient $2ln(BC)$.